Police response time to an emergency call is the difference between the time the call is first received by the dispatcher and the time a patrol car radios that it has arrived at the scene. Over a long period of time, it has been determined that the police response time has a normal distribution with a mean of 11.6 minutes and a standard deviation of 2.3 minutes. For a randomly received emergency call, find the following probabilities. (Round your answers to four decimal places.)

(a) the response time is between 7 and 13 minutes

(b) the response time is less than 7 minutes

(c) the response time is more than 13 minutes
Answer #1

TOPIC:Normal distribution.

→ Let ue dedine, x = The police response time (in minutes n so, / X ~ N(M = 11:6, 8 = 2.31 7 X- P 13 > -11.6 2.3 -11-6 2.3 @by we need, *P(x<7). - ? (*<?=0). =P ( 2 (-2. 7-11.6 2.3 D(-2). = 1- 2 (2). (From Normal = 1-0. 9772. = 10.0228 (Ans cy we n